### v1

# Multi-Level Explanations for Generative Language Models

Abstract
Perturbation-based explanation methods such as LIME and SHAP are commonly applied to text classification. This work focuses on their extension to generative language models. To address the challenges of text as output and long text inputs, we propose a general framework called MExGen that can be instantiated with different attribution algorithms. To handle text output, we introduce the notion of scalarizers for mapping text to real numbers and investigate multiple possibilities. To handle long inputs, we take a multi-level approach, proceeding from coarser levels of granularity to finer ones, and focus on algorithms with linear scaling in model queries. We conduct a systematic evaluation, both automated and human, of perturbation-based attribution methods for summarization and context-grounded question answering. The results show that our framework can provide more locally faithful explanations of generated outputs.
1 Introduction
Perturbation-based explanation methods, such as LIME [22] and SHAP [14], are popular for explaining individual predictions of black-box classification and regression models. The explanations take the form of attribution scores assigned to each part of the input, quantifying the effect of that part on the prediction when perturbed. [^1] We use the broader term “input attribution” in this paper rather than “feature attribution” because “features” may no longer be the best term for input parts such as text spans. For text input, the input parts are spans of text (often words), and applications to text classification tasks such as sentiment analysis have been well-studied [2, 13, 17, 12].
This work focuses on perturbation-based explanations for generative language tasks such as summarization and question answering. There has been little work on such explanations for text generation as corroborated by a survey paper on SHAP-based methods [17], the most relevant work being a collection of text examples in the SHAP package documentation [25] and not a paper. Concurrent with this work, the Captum library has added new capabilities for generative language models [16].
Perturbation-based input attribution for generative tasks presents challenges related to having text as output and long texts as input. We propose a general framework called MExGen (Multi-level Explanations for Generative Language Models) to address these challenges. The framework can be instantiated with different attribution algorithms, and we do so using SHAP- and LIME-like algorithms.
Figure 1: Example of using MExGen C-LIME (BERT scalarizer) to generate explanations. The example shows how the algorithm captures copied text (further skeletal remains) but also abstract concept mappings (M54 $\rightarrow$ motorway).
Figure 2: Diagram showing the workflow of MExGen.
Challenge of output text
The first challenge is that language models output text rather than a real number (e.g., the predicted log probability of a class). Attribution algorithms require a real-valued function to quantify that function’s sensitivity to different inputs. We address this through the concept of scalarizers, functions that map output text to real numbers. We investigate the use of multiple scalarizers and compare the resulting attributions.
Importantly, most scalarizers that we consider address the truly “black-box” setting in which we receive only text as output from the model. This is common with large language models (LLM) that only provide API access or are proprietary. In this case, scalarizers can use only the output text from the model being explained. The text-only setting is not handled by Captum and is only briefly alluded to in SHAP [25].
Challenge of input length
The second challenge is that the input text for generative tasks can be long, particularly for summarization and context-grounded question answering. Longer inputs require more resources, in terms of input perturbations and model queries, and impose financial costs for LLMs with API-only access. Long inputs also pose interpretation issues, since it is not clear what level of granularity in the attributions is most useful to a user. We address the challenge in three ways:
Linguistic segmentation. We segment the input text into linguistic units at multiple levels, for example sentences, phrases, and words, taking advantage of linguistic structure present in the text.
Multi-level attributions. We use a refinement strategy that proceeds from attributing at a coarser level like sentences to a finer level like phrases or words, with optional user input. This controls the number and granularity of units, and limits the compute.
Linear-complexity algorithms. We instantiate our framework with attribution methods whose number of perturbations and model queries scale linearly with the number of units at any level (e.g., number of sentences). In the SHAP category, we choose Local Shapley (L-Shapley) [4]. For LIME-type methods, we propose a linear-complexity variant that also limits the number of units perturbed at one time.
We evaluate the instantiations of MExGen on the tasks of summarization and context-based question answering. The automated evaluation shows that instantiations of MExGen are more locally faithful; particularly, they assign the highest importance to the input parts that are more relevant to the model output with higher frequency than the baselines. Automated evaluations also show that different scalarizers may yield similar explanations depending on the model being explained and the task being performed. Specifically, we found that some scalarizers — which only use API access to the model — generate explanations that closely resemble explanations using the log probabilities of the model being explained. Human evaluation results not only corroborated the automated evaluation findings but also provided additional insights into how certain scalarizers or attribution methods, previously considered similar to other methods in automatic evaluation, were perceived as more faithful by users.
Below is a summary of our main contributions:
  • We propose the MExGen framework to extend perturbation-based input attribution to generative language models, with a multi-level strategy to combat the challenges of long inputs.
  • We investigate several scalarizers for mapping output text to real numbers, notably handling the case of text-only output from the model.
  • We conduct a systematic evaluation, both automated and human, of input attribution methods for summarization and question answering, showing that MExGen can provide more locally faithful explanations of generated output than the alternative methods available.
2 Related Work
Post Hoc Explanations for LLMs
We focus on perturbation-based post hoc explanation methods for LLMs, notably those with only API access where text is the only available output.
Gradient-based methods provide input attribution explanations [27, 29, 26, 1], but they require access to model gradients with respect to the input.
While there is a significant literature on explaining text classification models, the literature on perturbation-based explanations for natural language generation is scarce, as noted in a survey of SHAP-based methods [17]. Below we discuss the two most relevant works that we have found. One common trait is that they require model logits (possibly from a proxy model) to generate explanations.
PartitionSHAP, the default algorithm for text in the SHAP library, handles long inputs by partitioning them into clusters of tokens and attributing the same score to tokens within a cluster. We have not found a paper describing PartitionSHAP. The description of the SHAP library’s text examples [25] is also cursory. However, it is apparent that PartitionSHAP produces one set of attributions for each token in the output. This approach is less interpretable because it requires the selection of an output token and assigns multiple attribution scores to each input span. PartitionSHAP can work with API-only access [24], but understanding how it does so required significant investigation on our part (see Appendix A.1).
CaptumLIME is our name for a modification of LIME [22] tailored for text generation tasks using new features in the Captum library [16]. [^2] Captum also has variants of SHAP but we found them slow to run and obviated by the availability of PartitionSHAP. CaptumLIME allows the user to manually define units for attribution within the input. It handles sequence outputs by computing log probabilities for tokens in a given output and then summing. This does, however, depend on access to output probabilities, so CaptumLIME is not suitable for the text-only setting.
Method LT Output LT Input API Access
LIME
SHAP
HEDGE
P-SHAP
Captum
MExGen
Table 1: Comparing the features (Long Text Output/Input, API access) of our MExGen framework with existing methods.
Hierarchical Explanations
A line of work [28, 11, 2, 12] has developed hierarchical explanations for sequence models, including LLMs, which can reveal compositional interactions between words and phrases. In particular, the HEDGE algorithm of Chen et al. [2] was identified by Mosca et al. [17] as “arguably the most suitable choice” for NLP input attribution, in part because it builds its hierarchy in a top-down, divisive fashion (as opposed to bottom-up agglomeration [28, 12]), which is more practical for long texts. However, HEDGE is specific to classification because it measures feature importance based on classification margin. Table 1 shows that PartitionSHAP, Captum, and HEDGE each lack capabilities compared to MExGen.
Natural Language Self-Explanations
We also note works have utilized the generative model itself to provide explanations in line with subsequent outputs, referred to as chain-of-thought or CoT [33]. These methods have been shown to be both unstable [32] and highly variable in quality [18]. Further, CoT provides self-explanations in natural language, and not input attribution explanations.
3 Multi-Level Explanations for Generative Language Models
We describe the proposed MExGen framework for input attribution for generative language models. Figure 2 provides an overview.
In the setting of perturbation-based input attribution, we are given a generative LLM $f$, an input text sequence of interest $x^{o}$ (left side of Figure 2, superscript $o$ for “original”), and a generated output $y^{o}=f(x^{o})$ that is also a text sequence (top middle of Figure 2). Our goal is to explain the output $y^{o}$ by attributing to parts of the input $x^{o}$. Each part of the input, denoted $x_{s}$, $s=1,\dots,d$, is to be assigned an attribution score $\xi_{s}$ (represented by color on the right of Figure 2) quantifying the importance of $x_{s}$ in generating the output, in the sense that if important parts are perturbed, then the output will change significantly. As the second through fourth paths in Figure 2 indicate, model $f$ can be queried on perturbations $x$ of $x^{o}$, with no further access.
Generative language tasks pose two main challenges: having text as output, and potentially long text as input. The following two subsections discuss our solutions to these challenges.
3.1 Handling Text Outputs
Input attribution methods require a numerical target function as the object to explain. Since an LLM $f$ may only output text, we introduce scalarizers, which are functions $S$ that map output text back to real numbers (shown as blue boxes in Figure 2). We consider two types of access to $f$: (a) full logit access, where $f$ can provide predicted logits for all tokens in its vocabulary, at each position in the output sequence, (b) text-only setting, where we are limited to text outputs. See Appendix A.1 for why vocabulary-wide access is assumed for (a).
Full logit access
When all logits are available, we use the probability of generating the original output sequence $y^{o}$ as the function to explain. We refer to this as the Log Prob scalarizer. Given output sequence $y^{o}$ of length $\ell$ and an arbitrary input sequence $x$, we compute from the model’s logits the log probability of generating each original output token $y^{o}_{t}$ conditioned on $x$ and previous output tokens $y^{o}_{<t}$. We then average over the output sequence to obtain the scalarized output for $x$,
$S(x;y^{o},f)=\frac{1}{\ell}\sum_{t=1}^{\ell}\log p\left(y^{o}_{t}\>|\>y^{o}_{<t},x;f\right).$ (1)
Here the scalarizer $S$ is parameterized by $y^{o}$ and $f$ since the latter is providing predicted probabilities.
The Log Prob scalarizer generalizes the log probability used in explaining text classification. This is seen from (1) by setting $\ell=1$ (single prediction) and identifying $y^{o}_{1}$ with the predicted class.
Text-only access
Using the output sequence $y=f(x)$ generated from the input $x$, we consider similarity measures $S(y;y^{o})$ between $y$ and the original output $y^{o}$ as scalarizers. These now depend on $f$ only via composition, i.e., $S(f(x);y^{o})$, and are not parameterized by $f$ as in (1).
  • Sim: Computes the cosine similarity between embedding vectors for $y$ and $y^{o}$.
  • BERT: Computes the BERTScore [35] between $y$ and $y^{o}$.
[…]
  • Log NLI: Uses a natural language inference (NLI) model to predict the log-odds of entailment given $y^{o}$ as premise and $y$ as hypothesis, and optionally in the other direction as well (see Appendix A.1).
3.2 Handling Long Text Inputs
Some generative language tasks require long input texts, such as in summarization and context-grounded QA. We address this challenge through a combination of three techniques: segmenting the input into linguistic units at multiple levels, using attribution algorithms with linear complexity in the number of input units, and obtaining attributions to units in a coarse-to-fine manner.
Linguistic segmentation
We segment the input text into linguistic units at multiple granularities: paragraphs, sentences, phrases, and words (“Select Units” box in Figure 2). This approach takes advantage of linguistic and other structure present in the input. For example, the input may already be broken into paragraphs or contain multiple distinct retrieved documents, in which case these paragraphs or documents can form the units at the highest level. In contrast, many existing methods rely on the model’s tokenizer, which can yield units (tokens) that are too fine, or treat the text as a flat sequence of tokens and let the algorithm decide how to segment it [2, 25].
We use tokenization and dependency parsing from spaCy v3.6 [9] to segment paragraphs into sentences and words. To segment sentences into phrases, we propose an algorithm that uses the dependency parse tree from spaCy. In the first pass, the algorithm recursively segments the tree and its subtrees into phrases that are no longer than a maximum phrase length. In the second pass, some short phrases are re-merged. More details are in Appendix A.2.
Our framework allows for units at any level to be marked as not of interest for attribution. For example at the word or token level, punctuation and stop words may be ignored. In question answering, prompts for LLMs often follow a template, e.g. prefacing the context with the string “Context: ”. Such elements of the template are usually not of interest. We may also not attribute to the question and focus only on the context.
Linear-complexity algorithms
Given an input segmentation (which may have mixed units as in Figure 2), we have the actual task of attributing to units $x_{1},\dots,x_{d}$ (the “Post Hoc Explainer” block in Figure 2). For this we consider only attribution algorithms that scale linearly with the number of units $d$ in terms of model queries, to control this cost. We instantiate MExGen with three such algorithms: Local Shapley (L-SHAP), a LIME-like algorithm with further constraints (C-LIME), and leave-one-out (LOO) as a baseline.
L-SHAP: This local variation of SHAP was proposed by Chen et al. [4] to decrease the number of model inferences relative to SHAP, which requires exponentially many inferences. In our context, L-SHAP does so by only perturbing units that are within a constant-size neighborhood of the current unit being attributed to. This makes the number of inferences scale linearly with the number of units. More precisely, for unit of interest $s\in[d]$, we consider only the radius-$M$ neighborhood $s-M,\dots,s+M$ (truncated to $1,\dots,d$ if necessary), and define $\Omega^{M,K}_{s}$ to be the set of all subsets of $s-M,\dots,s+M$ up to cardinality $K$. Then the attribution score $\xi_{s}$ for unit $s$ is given by
$\xi_{s}$ $=\sum_{A\in\Omega^{M,K}_{s}}\frac{S(x^{(A)};y^{o},f)-S(x^{(A\cup\{s\})};y^{o},f)}{Z_{|A|}},$ (2)
where $x^{(A)}$ is a perturbation of $x^{o}$ in which units $j\in A$ are replaced (e.g., by a baseline value such as [MASK] or using a LM), and $Z_{|A|}$ is the number of subsets in $\Omega^{M,K}_{i}$ with cardinality $|A|$.
C-LIME: Similar to LIME, we use a linear model that operates on interpretable features $z$ and approximates the model $f$ in the vicinity of original input $x^{o}$. In our case, the interpretable features $z\in\mathbb{R}^{d}$ correspond to the units $x_{1},\dots,x_{d}$, with $z_{s}=0$ if unit $s$ is perturbed and $z_{s}=1$ otherwise. The linear model is fit using perturbations $x^{(1)},\dots,x^{(n)}$ of $x^{o}$, with corresponding interpretable representations $z^{(1)},\dots,z^{(n)}$, and model outputs (scalarized in our case) $S(x^{(1)};y^{o},f),\dots,S(x^{(n)};y^{o},f)$:
$\xi=\operatorname*{arg\,min}_{w}\sum_{i=1}^{n}\pi\bigl{(}z^{(i)}\bigr{)}\bigl{(}w^{T}z^{(i)}-S(x^{(i)};y^{o},f)\bigl{)}^{2}+\lambda R(w),$ (3)
where $\pi(z)$ is a sample weighting function and $R$ is a regularizer. The best-fit linear model coefficients yield the attribution scores $\xi$.
We make two main departures from LIME, in addition to the use of a general scalarizer $S$ as seen in (3). Firstly, we limit the number of perturbations $n$ to a multiple of the number of units $d$. Since the regression parameters $w$ are $d$-dimensional (not including a possible intercept), the multiplier could be set to $10$ for example to have $10$ times as many samples (i.e., perturbations) as parameters to fit. In contrast, LIME by default sets $n$ to be in the thousands independently of $d$, which can be prohibitive for LLMs. Secondly, we limit the number of units $K$ that can be perturbed simultaneously to a small integer. This concentrates the smaller number of perturbations on inputs that are closer to $x^{o}$, which has been shown to improve the local fidelity of attributions [30]. LIME by contrast samples the number of units to perturb uniformly from $\{1,\dots,d\}$ [15]. Other details are in Appendix A.3.
LOO: We also consider a LOO variant of MExGen as a baseline. Here, units $x_{1},\dots,x_{d}$ are perturbed one at a time to yield perturbations $x^{(1)},\dots,x^{(d)}$ (i.e., $n=d$). The attribution score for $x_{s}$ is the corresponding decrease in scalarizer score from its original value, $\xi_{s}=S(x^{o};y^{o},f)-S(x^{(s)};y^{o},f)$.
Multi-level explanations
Once we have computed attribution scores at a given level, we may choose to refine the input units and repeat the process (feedback path at the bottom of Figure 2). For example, given sentence-level attributions for a long input, we may wish to obtain phrase- or word-level attributions for a few sentences and keep attributions at the sentence level for the remainder to not introduce too many new units. The few sentences can be selected by any number of automatic criteria or by the user, or the user can even choose to stop at sentence level. See Appendix B for sentence selection criteria used in our experiments.
4 Automated Evaluation
We evaluate MExGen across two text generation tasks: Context-grounded QA and summarization.
4.1 Setup
Datasets
For summarization, we use two popular datasets, Extreme Summarization (XSUM) [19] and CNN/Daily Mail (CNN/DM) [23, 8]. For question answering, we use the Stanford Question Answering Dataset (SQuAD) [21] because of its relatively short, paragraph-long contexts, which allow us to test mixed sentence-and-word-level attributions.
Models
For summarization, we use two models, a 306M-parameter distilbart-xsum-12-6 model [^3] https://huggingface.co/sshleifer/distilbart-xsum-12-6, and a larger 20B-parameter Flan-UL2 model [31]. We treat the DistilBart model as one with full logit access, thus enabling use of the Log Prob scalarizer. We call Flan-UL2 through an API. [^4] We use Flan-UL2 via the IBM Generative AI Python SDK (Tech Preview) [10]. For QA, we use the 770M-parameter Flan-T5-Large model [^5] https://huggingface.co/google/flan-t5-large [5] and treat it as providing output logits.
Attribution algorithms
We instantiate MExGen with the attribution algorithms discussed in Section 3.2: L-SHAP, C-LIME, and LOO, where the third is intended more as a baseline. For summarization, we obtain mixed sentence and phrase attributions while for QA, we obtain mixed sentence and word attributions. Algorithm parameter settings are given in Appendix B. We compare against PartitionSHAP (P-SHAP) [25] and CaptumLIME [16]. For PartitionSHAP, recall from Section 2 that it requires the selection of an output token to explain, whereas we are interested in the output sequence as a whole. For this reason, we slightly modify PartitionSHAP by summing across all attribution scores, corresponding to different output tokens, that it gives to each input span. This is equivalent to explaining the sum of the log probabilities of output tokens because of the linearity of Shapley values, and it corresponds to our Log Prob scalarizer. For CaptumLIME, since it accepts user-defined units for attribution, we provide it with the same sentence-level and mixed-level units used by our algorithms. Thus we can directly compare Captum’s attribution algorithm (i.e., LIME) with ours, controlling for input segmentation. Captum’s default for the target function to explain also corresponds to the Log Prob scalarizer. Additionally, PartitionSHAP and CaptumLIME take the number of model queries as an input; for a fair comparison, we allow them the greater of the numbers of model calls used by MExGen L-SHAP and C-LIME.
(a) MExGen C-LIME
(b) MExGen L-SHAP
Figure 3: Spearman’s rank correlation between attribution scores using different scalarizers. Attributions were computed using multi-level (a) MExGen C-LIME for the DistilBart model on the XSUM dataset and (b) MExGen L-SHAP for Flan-T5-Large on the SQuAD dataset.
Metrics
To measure local fidelity of explanations, we use perturbation curves as in previous work on explaining text classification [2, 12]. The general idea behind perturbation curves is to rank input units in decreasing importance according to a set of attribution scores, perturb the top $k$ units with $k$ increasing, and plot the resulting change in some output score that measures how much the perturbed output deviates from the original output. For the output score, we select an evaluation scalarizer from the set of scalarizers in Section 3.1. To accommodate PartitionSHAP and CaptumLIME, we choose an evaluation scalarizer corresponding to the target function that they use for explanation. For the DistilBart summarization and Flan-T5-Large QA models, this means the Log Prob scalarizer, while for Flan-UL2, we choose SUMM in keeping with PartitionSHAP. Appendix C has further details on how we compute perturbation curves, accounting for the different units used by different attribution algorithms.
4.2 Scalarizer Evaluation
Unit Ranking Similarity Across Scalarizers
We evaluate the similarity between attribution score vectors $\xi_{S}$ and $\xi_{S^{\prime}}$ at the same granularity level for all pairs of scalarizers $S,S^{\prime}$. We present Spearman rank correlation to measure similarity in unit rankings. See Appendix C for further results.
Figure 3 shows Spearman correlation matrices, averaged across examples from the respective datasets. Certain pairs are highly similar, for example BART and SUMM as mentioned in Section 3.1, as well as Sim and BERT. Log Prob, the only one that uses logits from the model being explained, clearly differs from the others. There are also noticeable differences between panels (a) and (b) in Figure 3, which feature different tasks, datasets, and models. For example, Spearman correlation between BERT and SUMM is $0.75$ in (a) and $0.83$ in (b). This suggests the necessity of exploring different scalarizers to explain different tasks.
(a) BERT Score as Evaluation
(b) Log Prob as Evaluation
(c) BART Score as Evaluation
Figure 4: Perturbation curves for MExGen C-LIME with different scalarizers, used to explain the DistilBart model on the XSUM dataset. The curves show the decrease in (a) BERTScore, (b) log probability, and (c) BARTScore when removing the most important $p$of tokens according to each explanation scalarizer. Shading shows standard error.
Perturbation Curves Across Scalarizers
Figure 4 shows perturbation curves with different evaluation scalarizers for MExGen C-LIME using different scalarizers to compute attribution scores (i.e., a cross-evaluation of scalarizers). Results are again averaged over examples from the datasets (see Appendix C for how). In panels (b) and (c), the case where the attribution scalarizer is matched with the evaluation scalarizer is generally the best. Additionally, the Log Prob scalarizer performs well across all three evaluation scalarizers. This implies that the Log Prob scalarizer gives the most universal (least dependent on the evaluation scalarizer) explanation; hence, when log probabilities are available, we suggest using them.
In the case where model logits are not available (e.g., when only API access is given), the choice of scalarizer is not clear. However, if (hypothetically) one were to use Log Prob as the evaluation scalarizer, our results in Figures 3 and 4(b) indicate that BERT score is the scalarization method that best approximates explanations with Log Prob as the scalarizer. We include results for other tasks, models, and evaluation scalarizers in Appendix C.
4.3 Comparison Between Explainers
(a) Flan-UL2 on CNN/DM
(b) DistilBart on XSUM
(c) Flan-T5-Large on SQuAD
Figure 5: Perturbation curves for different explanation methods, using the same scalarizer as the evaluation indicated on the y-axis. Models and datasets: (a) Flan-UL2 on the CNN/DM dataset, (b) DistilBart on XSUM, and (c) Flan-T5-Large on SQuAD. MExGen provides multi-level explanations down to phrase-level for XSUM and CNN/DM and word-level for SQuAD. We also show the perturbation curve of MExGen C-LIME using mismatched scalarizers: BARTScore (a) and BERTScore (b)(c). Shading shows standard error in the mean.
Perturbation Curves
We compare the performance of the MExGen instantiations with P-SHAP and CaptumLIME. We were only able to generate results for CaptumLIME in Figure 5 (b) because (i) we had API access to Flan-UL2 and CaptumLIME needs output logits, and (ii) CaptumLIME does not support Flan-T5-Large. Figure 5 shows that perturbing units (i.e., the tokens that they comprise) ranked as most important by MExGen has the largest influence on the evaluation scalarization, compared to P-SHAP and CaptumLIME. This indicates that the methods using MExGen identified more important tokens according to the evaluation scalarizer, while P-SHAP and CaptumLIME give higher scores for less important tokens.
We found that MExGen C-LIME consistently outperforms the rest of the methods across all tested models and datasets. In particular, the curve for MExGen C-LIME is always higher than that of P-SHAP up to at least the top 10% of tokens; the gray vertical lines in Figure 5 show the fraction of tokens where the cross-over occurs. We also show in Figure 5 the perturbation curve for MExGen C-LIME using a different scalarizer than used for evaluation. Surprisingly, we find that even when using a mismatched scalarizer (notably BERT in (b)(c), which does not even use logits), MExGen C-LIME can outperform P-SHAP in local fidelity.
Datasets Models MExGen C-LIME MExGen L-SHAP MExGen LOO P-SHAP
XSUM DistilBart
14.7 14.7 14.1 10.7
Flan-UL2 21.3 21.2 18.4 20.0
CNN/DM DistilBart
13.5 14.7 13.2 9.7
Flan-UL2 31.7 30.8 28.3 30.7
SQUaD Flan-T5-Large 61.2 59.9 59.9 59.4
Table 2: Area under the perturbation curve (AUPC, higher is better) up to $20$% of tokens. For DistilBart and Flan-T5-Large, log probability is used as both the explanation and evaluation scalarizer. For Flan-UL2, which does not provide access to log probabilites, SUMM is used as the explanation and evaluation scalarizer.
Area Under the Perturbation Curve
Table 2 shows the area under the perturbation curve (AUPC) to summarize performance over all dataset-model pairs that we tested, not just those in Figure 5. We evaluate AUPC up to $p=20\%$ as done in Chen et al. [2]. Across all dataset-model pairs, methods that use the MExGen framework (sometimes including the baseline LOO) performed better (i.e., higher AUPC) than P-SHAP. Moreover, the second highest AUPC is always a MExGen method.
5 User Study
We conducted a user study to understand how humans perceive explanations provided by different scalarizers and attribution methods, and whether they can discern performance differences akin to the quantitative evaluations in Section 4. To ease interpretation, we developed a visualization tool that highlights input text spans based on a color-coded scale for the attribution scores. We addressed the following research questions for summarization:
  1. 1.
    Fidelity: Which method is perceived to be better at explaining how the language model generated the summary?
  2. 2.
    Preference: Which method do people prefer?
  3. 3.
    Concentration: What are people’s perceptions regarding the concentration of attribution scores (i.e., do the darker colors cover less of the text)?
  4. 4.
    Granularity: Which granularity level (sentence-level vs. multi-level) do people prefer?
We selected ten examples from the XSUM dataset for the user study with diversity in topics, while ensuring that they do not contain sensitive issues or obvious errors. We designed an online survey consisting of three parts. Each part showed an input text, randomly drawn from the ten examples, and its summary, generated by the DistilBart model. This user study focused on algorithms and scalarizers that are of most interest based on the automated evaluation in Section 4. Specifically, we compared two scalarizers (Log Prob, BERT) in Part 1, three attribution algorithms (C-LIME, L-SHAP, PartitionSHAP) in Part 2, and two levels of granularity (sentence-level, multi-level) in Part 3. The presentation order of the attribution algorithms was randomized to mitigate order effects. The survey consisted of seven pairwise comparisons in total followed by questions for the participants.
We recruited participants from a large technology company who self-identify as machine learning practitioners using language models and collected data from 88 of them after filtering. Here, we report a summary of key results only. See Appendix D for details including survey questions, analysis, and statistical results.
Scalarizers. Significantly more participants perceived BERT to be higher in fidelity than Log Prob. They also preferred BERT over Log Prob. Participants perceived that the attribution concentration was adequate in both BERT and Log Prob.
Attribution methods. Significantly more participants perceived C-LIME to be higher in fidelity than L-SHAP. They also preferred C-LIME over L-SHAP. Participants perceived that the attribution concentration was adequate in all three methods.
Granularity. Regarding user preference between sentence-level and multi-level granularities, no statistically significant difference was found.
6 Concluding Remarks
We proposed MExGen, a framework to provide explanations for generative language models. Our experiments show that instances of MExGen provide faithful explanations and outperform the baselines we compare against — PartitionSHAP and CaptumLIME. Our numerical results also indicate that C-LIME outperforms the other explanation methods in the MExGen framework and PartitionSHAP. The user study results align with the automated evaluation overall, and additionally reveal that people perceive the BERT scalarizer as more locally faithful than the Log Prob scalarizer, which was not seen in the automated evaluation. This result implies that in some cases, there may be no loss in having text-only access compared to full logit access.
Limitations
We see three main limitations in our work. First, MExGen is a framework for post hoc explanations. Although providing such explanations can be helpful for practitioners to understand the model behavior, it does not fully characterize how models make predictions and only provides local explanations. Second, our automatic evaluations use three models (DistilBart, Flan-UL2, and Flan-T5-Large) and three datasets (XSUM, CNN/DM, and SQUaD). Although our findings are consistent across the models and datasets, the results reported in Section 4 could still change in different experimental settings. Third, our user study analyzes the perception of participants of how well a method explains the predictions of a given model, and not necessarily the fidelity of the explanation itself — fidelity is measured more directly in the automated evaluation. However, we believe the fact that the participant pool was composed of people with experience in ML and LLMs improves the quality of evaluation of the explanation methods. Finally, with post hoc explanations, there is the risk that the explanation algorithm could be steered to obfuscate undesirable behavior from the model.
Ethics Statement
MExGen is a framework to explain generative language models. Hence, its objective is to elucidate how a model made a specific prediction. Methods that aim to understand how black-box models generate their output are essential for guaranteeing transparency during decision-making. For example, a generative language model can be used to summarize dialog and create minutes of meetings that can later be used to perform high-stakes decisions. Then, it is necessary to understand how the model generated the summary and ensure that the output content is based on the input dialog. Therefore, for such high-stakes applications, methods that can provide explanations for text generated by language models are necessary, highlighting the importance of MExGen.
Acknowledgements
The authors thank the following people: Inkit Padhi for suggestions on parsing sentences into phrases, using language models for perturbing words, and participation in multiple discussions; Ella Rabinovich and Samuel Ackerman for discussions on and recommendations of similarity measures as scalarizers, and Ella in particular for also providing suggestions on parsing sentences into phrases and using language models for perturbation; Pratap Kishore Varma Vemulamanda for helping to investigate the API-only case of SHAP’s summarization example [24]; Subhajit Chaudhury for suggestions of similarity measures as scalarizers; Keshav Ramji for discussion on related work and quality evaluation for the explanations; Kush Varshney and Eitan Farchi for general encouragement and support.
Appendix A Further Details on MExGen
A.1 Scalarizers
Vocabulary-wide logits
The inputs $x$ in our case are perturbed versions of the original input $x^{o}$. Some of these perturbations however may be significantly different semantically and cause the probability of generating the original output $y^{o}$ to decrease dramatically. For this reason, the Log Prob scalarizer may require access to logits for improbable tokens conditioned on $x$, which can be ensured if logits are available for the entire vocabulary, but not if only the top $k$ logits are provided.
Aggregation for Log Prob scalarizer
As alternatives to the average over the output sequence in (1), other ways of aggregating include using the sum or taking the product or geometric mean of the probabilities. We choose the average to normalize for sequence length and because log probabilities tend to be a more “linear” function of inputs than probabilities.
BERTScore
BERTScore uses an LM to obtain contextual embeddings for the tokens in $y$ and $y^{o}$, matches the two sets of embeddings based on cosine similarity, and computes a score to quantify the degree of match.
SUMM scalarizer
[…]
Log NLI scalarizer
This scalarizer is based on the intuition that Log NLI entailment is a kind of similarity. We use the NLI model to predict the log-odds of entailment given $y^{o}$ as premise and $y$ as hypothesis, and optionally in the other direction as well. If both directions are used, we take the geometric mean of the two entailment probabilities and then convert back to log-odds.
A.2 Phrase Segmentation
Here we describe the phrase segmentation algorithm mentioned in Section 3.2. The algorithm starts with the root token of the tree and checks whether each child subtree of the root is shorter (in terms of tokens) than a maximum phrase length parameter. If it is shorter, then the subtree constitutes a phrase, and if it is not, the algorithm is recursively applied to the subtree. The root of each (sub)tree is also taken to be a phrase. Once the sentence has been recursively segmented into phrases in this manner, a second pass is performed to re-merge some phrases that have become too fragmented, thus controlling the number of phrases which is desirable for computation and interpretation. Specifically, phrases that constitute noun chunks (as identified by spaCy) are merged, and certain single-token phrases are merged with their neighbors. Further notes:
  • Subtrees of the dependency parse tree are usually contiguous spans of text (in English), but sometimes they correspond to multiple spans. In this case, we treat each span as its own subtree since we wish to have contiguous phrases.
  • In measuring the token length of a span, we do not count punctuation or spaces.
  • In merging phrases that fall within a noun chunk, we check conditions that are consistent with noun chunks, for example that there is a root phrase that is a single token (the noun), and that the other phrases are children of the root phrase.
  • For merging single-token phrases with their neighbors, we use the following criteria:
    • The single-token phrase (singleton) is a non-leaf phrase (is the parent of other phrases) or a coordinating conjunction.
    • If the singleton is a coordinating conjunction (e.g. “and”), the neighbor is a corresponding conjunct (e.g. “Bob” in “Alice and Bob”).
    • If the singleton is a preposition (e.g. “to”), the neighbor is a child of the preposition (e.g. “the store” in “to the store”).
    • If the singleton is of some other type, the neighbor is a leaf phrase and is either adjacent to the singleton or a singleton itself.
    • The merged phrases do not exceed the maximum phrase length parameter.
A.3 C-LIME
In addition to the departures from LIME described in Section 3.2, we discuss two other aspects:
Sample weighting
Since we limit the number of units that are simultaneously perturbed to a small integer $K$, we also no longer use LIME’s sample weighting scheme. Instead, we give each subset cardinality $k=0,\dots,K$ the same total weight, and then distribute this weight uniformly over the subsets of that cardinality.
Regularization
Different regularizers $R(w)$ can be used in (3), e.g. $\ell_{2}$ or $\ell_{1}$. In our experiments however, we do not regularize and compute a fully dense solution. This allows all units to be ranked to facilitate evaluation of perturbation curves.
Appendix B Algorithm Parameter Settings
This appendix documents the settings of key algorithm parameters and other choices made in the experiments.
Scalarizer models
The text-only scalarizers presented in Section 3.1 can be instantiated with different models. The ones used in our experiments are as follows:
  • Sim”: We use the all-MiniLM-L6-v2 embedding model from the SentenceTransformers package666this model can be found at https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2.
  • BERT”: We use the Log NLI model deberta-v2-xxlarge-mnli777this model can be found at https://huggingface.co/microsoft/deberta-v2-xxlarge-mnli [7] to compute BERTScore, which is the same Log NLI model that we use for the Log NLI scalarizer. Our initial reason for doing so was to see whether the two scalarizers would be very similar because of this choice (they are not as they operate on different principles). We take the F1-score output as the BERTScore.
  • BART” and “SUMM”: For “SUMM”, we follow SHAP [24] in using the same distilbart-xsum-12-6 model as both the scalarizing summarization model as well as the primary summarization model to explain. For “BART”, we also use distilbart-xsum-12-6 to determine whether the “BART” and “SUMM” scalarizers are very similar when instantiated with the same model. This is indeed the case.
  • Log NLI”: As mentioned above, we use deberta-v2-xxlarge-mnli as the Log NLI model. We also choose to compute the Log NLI entailment probability in both directions and take the geometric mean of the two before taking the logit.
All other parameters of these scalarizers are kept at default values.
Linguistic segmentation
The phrase segmentation algorithm has one main parameter, the maximum phrase length, which we set to $10$ spaCy tokens (not counting spaces and punctuation). We exclude non-alphanumeric units (generally punctuation) from attribution. For QA, we allow only the context to be attributed to.
Multi-level explanations
The decision here is how many top sentences are refined into phrases in the summarization experiments and into words in the QA experiments. For summarization, this is determined by first normalizing the sentence-level attribution scores to the range $[-1,1]$ and then choosing sentences with absolute attribution scores above a threshold. The threshold is $1/3$ for C-LIME and LOO with the distilbart-xsum-12-6 model, $0.3$ for L-SHAP with distilbart-xsum-12-6, $0.5$ for C-LIME and LOO with the Flan-UL2 model, and $0.4$ for L-SHAP with Flan-UL2. If more than three sentences have scores above the threshold, we take only the top three. For QA, we simply take the top $1$ sentence as the context paragraphs in SQuAD tend to have only a handful of sentences.
L-SHAP
For L-SHAP, the two main parameters are the local neighborhood radius $M$, which we take to be $M=2$, and the maximum number of units $K$ perturbed at one time, also set to $2$. Note that the latter does not include the unit of interest, so altogether the maximum is $3$.
C-LIME
For C-LIME, the main parameters controlling the perturbations are the constant of proportionality between the number of perturbations $n$ and number of units $d$, and the maximum number of units $K$ perturbed at one time. The former is set to $10$ and the latter to $3$, except for sentence-level attributions on SQuAD where it is $2$.
PartitionSHAP
We set the number of model queries (parameter max_evals) to be approximately equal to the number used by our L-SHAP and C-LIME algorithms. More specifically, since obtaining mixed-level attributions with MExGen requires first performing sentence-level attribution, we add the numbers of queries used during sentence-level and mixed-level attribution. We then take the larger of these two sums for L-SHAP and C-LIME as the number of queries allowed for PartitionSHAP.
Appendix C Automated Evaluation Details and More Quantitative Results
Computing environment
Experiments were run on a computing cluster providing V100 and A100 GPUs with 32 GB of CPU memory and 32, 40, or occasionally 80 GB of GPU memory. One CPU and one GPU were used at a time. The total computation time is estimated to be on the order of 100–200 hours.
Datasets
For summarization, we limit our evaluation to the first 100 examples from the test sets of XSUM and CNN/Daily Mail, due to the computation involved in running multiple perturbation-based attribution algorithms that query summarization models. For question answering on SQuAD, we evaluate 200 randomly sampled examples from the validation set.
Perturbation curves
MExGen can attribute to mixed units of different lengths in terms of the number of tokens, and these units also differ from those produced by P-SHAP. To account for these differences in computing perturbation curves, we consider both the attribution score and the number of tokens for each unit. We rank units in decreasing order of the unit attribution score divided by the number of tokens (since we plot perturbation curves as functions of the number of tokens perturbed, this ratio can be seen as the slope in the score-tokens plane). We then perturb (more precisely remove) the top $k$ units according to these ratios, compute the output score given by the evaluation scalarizer, and increase $k$ until at least $50$% of the tokens have been removed. To average perturbation curves over examples with different numbers of tokens, we divide the numbers of tokens perturbed by the total number to obtain percentages. We then linearly interpolate onto a common grid of percentages before averaging.
(a) (C-LIME, DistilBart, XSUM)
(b) (L-SHAP, DistilBart, XSUM)
(c) (C-LIME, DistilBart, XSUM)
(d) (L-SHAP, DistilBart, XSUM)
(e) (C-LIME, Flan-T5, SQUaD)
(f) (L-SHAP, Flan-T5, SQUaD)
(g) (C-LIME, Flan-T5, SQUaD)
(h) (L-SHAP, Flan-T5, SQUaD)
Figure 6: Spearman’s rank correlation and cosine similarity for different explanation methods in varying datasets.
On the Cosine Similarity of Explanations
Figure 6 shows the cosine similarity between all the pairs of secalarizers we use. We show the cosine similarity between the explanations of each method to analyze how aligned the explanations are from different scalarizers. Each text input $x=x_{1},...,x_{d}$ receives a multi-level explanation given by $\xi_{S}(x)=(\xi_{S}(x)_{1},...,\xi_{S}(x)_{d})\in\mathbb{R}^{d}$ where each $\xi_{S}(x)_{i}$ represents the contribution of unit $i$ to the model prediction scalarization computed using the scalarizer $S$. We define the cosine similarity between scalarizers $S$ and $S^{\prime}$ as the average of the cosine similarities between the explanations for all available input texts, i.e.,
$\texttt{CosSim}(S,S^{\prime})\triangleq\frac{1}{|X|}\sum_{x\in X}\frac{\langle\xi_{S}(x),\xi_{S^{\prime}}(x)\rangle}{||\xi_{S}(x)||||\xi_{S^{\prime}}(x)||}.$
Figure 6 (a) shows the cosine similarities for the scalarizers used by MExGen C-Lime to explain the predictions of the DistilBert model in the XSUM dataset. Figure 6 (b) shows the cosine similarities for the scalarizers used by MExGen L-SHAP to explain the predictions of the Flan-T5-Large model on the SQUaD dataset.
[…]
Similarity to Logist. In the SHAP library, SUMM is proposed to provide explanations to LLMs that do not provide access to the logits — hence, the main objective of SUMM is to approximate the explanations for the logit when it is not available. However, Figure 6 shows that SUMM is not always the best scalarizer for approximating the explanations that would be given if logits were available. For example, Figure 6 (a) shows that the similarity between SUMM and logit is near $0.61$. In contrast, Figure 6 (b) shows that the similarity between the explanations generated using SUMM as scalarizer has a similarity of $0.78$ with the onex generated using logits.
We are also aware that only comparing the similarities between explanations might not be enough; once, many researchers use the scores to compute the ranking of the importance across all input features (input text units here). For this reason, next, we compare Spearman’s rank correlation to measure the rank stability across different scalarizers.
Similarity to Log Probability Scalarizer. In the SHAP library, SUMM is proposed to provide explanations when access to logits is not available. However, Figure 3 shows that SUMM is not always the best scalarizer for approximating the explanations that would be given if logits were available. For example, Figure 3 (a) shows the ranking correlation between the scalarizer and the Log Prob scalarizer is higher for BERT score. Figure 3 (b) shows that the rank generated by both the explanations generated using SUMM and BERT scalarizer are equally similar to the rank of explanations using the Log Prob scalarizer.
Perturbation curves for different combinations of scalarizers
Figures 7, 8, 9 show the perturbation curve for different scalarizations.
(a) BERT Score as Evaluation
(b) Log Prob as Evaluation
(c) BART Score as Evaluation
Figure 7: Perturbation curves for MExGen L-SHAP with different scalarizers, used to explain the DistilBart model on the XSUM dataset. The curves show the decrease in (a) BERTScore, (b) log probability, and (c) BARTScore when removing the most important $p$of tokens according to each explanation scalarizer. Shading shows standard error.
(a) BERT Score as Evaluation
(b) Log Prob as Evaluation
(c) BART Score as Evaluation
Figure 8: Perturbation curves for MExGen C-LIME with different scalarizers, used to explain the Flan-T5-Large model on the SQUaD dataset. The curves show the decrease in (a) BERTScore, (b) log probability, and (c) BARTScore when removing the most important $p$of tokens according to each explanation scalarizer. Shading shows standard error.
(a) BERT Score as Evaluation
(b) Log Prob as Evaluation
(c) BART Score as Evaluation
Figure 9: Perturbation curves for MExGen L-SHAP with different scalarizers, used to explain the Flan-T5-Large model on the SQUaD dataset. The curves show the decrease in (a) BERTScore, (b) log probability, and (c) BARTScore when removing the most important $p$of tokens according to each explanation scalarizer. Shading shows standard error.
Appendix D User Study
D.1 Participants
We recruited 96 participants from a large technology company. Those who self-identified as machine learning practitioners using language models were eligible for the study. We filtered out 8 participants who did not pass eligibility checks or did not provide valid responses, resulting in data from 88 participants for our analysis.
D.2 Scalarizers
[…]
Selected Option C-LIME L-SHAP
Log Prob 35.2% 34.1%
BERT 54.5% 60.2%
Identical 10.2% 5.7%
Table 3: The proportions of participants who selected one of the three options – Log Prob, BERT, or ‘They are identical’. Regardless of attribution methods, significantly more participants chose BERT over Log Prob when asked to select the one with higher perceived fidelity.
Preferred Option C-LIME L-SHAP
Log Prob 29.5% 31.8%
BERT 62.5% 64.8%
Identical 8% 3.4%
Table 4: The proportions of participants who selected one of the three options – Log Prob, BERT, or ‘They are identical’ based on their preference. Regardless of attribution methods, significantly more participants preferred BERT over Log Prob.
D.3 Attribution Methods
[…]
Selected vs. Rejected Options $p$-value
C-LIME vs. L-SHAP 0.0107 **
PartitionSHAP vs. L-SHAP 0.0455
C-LIME vs. PartitionSHAP 0.5691
Table 5: There is a significant difference in perceived fidelity between C-LIME and L-SHAP. Significant p-values after Bonferroni adjustment are noted with ** (p<0.05/3).
Selected vs. Rejected Options $p$-value
C-LIME vs. L-SHAP 0.0074 **
PartitionSHAP vs. L-SHAP 0.1758
C-LIME vs. PartitionSHAP 0.1758
Table 6: There is a significant difference in preference between C-LIME and L-SHAP. Significant p-values after Bonferroni adjustment are noted with ** (p<0.05/3).
D.4 Granularity Preference
Participants were asked to select their preferred granularity of attributions (sentence-level vs. multi-level). While the number of participants who preferred multi-level granularity (56.2%) was slightly higher than those who preferred sentence-level granularity (43.8%), binomial tests indicated that their granularity choice was not statistically significant. The preference for granularity did not significantly vary across attribution algorithms (C-LIME, L-SHAP).
D.5 Survey
Figure 10 and Figure 11 show some questions we asked in the survey.
Figure 10: Survey questions. Participants answered a series of questions related to perceived fidelity and general preference comparing a pair of attribution methods, followed by questions related to perceived concentrations of each method.
Figure 11: Granularity question. Participants were asked to select their preferred granularity of attributions (sentence-level vs. multi-level).
Appendix E Future Directions
Hierarchical explanations
It could be profitable in future work to incorporate the hierarchical explanations discussed in Section 2 into the multi-level MExGen framework. The method of Chen and Jordan [3] may be especially relevant since it leverages a constituency parse tree to compute word-level importances, which may be related to our use of dependency parse trees.
Word infilling with BERT
We have explored perturbing words by masking them and then calling a BERT model to fill the masks with different words that fit within the sentence. However, we have thus far not seen a quantifiable benefit to using BERT compared to replacing with a fixed baseline value (such as an empty string). Our experience is in line with the the mixed results reported by Pham et al. [20] on using BERT in this manner.
Phrase segmentation
Segmentation of sentences into phrases could of course be done in ways other than our dependency parsing algorithm, for example using constituency parsing instead. A possible advantage of using dependency parsing is that each phrase can be labeled with the dependency label of its root token and treated differently on this basis.